OpenStudy (anonymous):

what is the slope of the line containing the pair of points (5,4) and (7,-6)

OpenStudy (anonymous):

would the answer be -5

OpenStudy (anonymous):

Yes it would

OpenStudy (anonymous):

Use the formula. (y2-y1)/(x2-x1). You get (-6-4)/(7-5) = -10/2 = -5